TI need your help, please...
The Product of two consecutive numbers is twice the square of the smaller number. Find the numbers.

Answer by stanbon(75887)

The Product of two consecutive numbers is twice the square of the smaller number. Find the numbers.
---
1st: x
2nd: x+1
-----
Equation:
x(x+1) = 2x
x^2+x-2x = 0
x^2-x = 0
x(x-1) = 0
---
Solutions:
x = 0 or x = 1
======================
Cheers,
Stan H.